Find the surface area
a. cylinder and cone 6mm, h=10mm, r=4mm
b. cone r=10mm, h=25mm
c. square base 50mm, 40mm, 40mm

Answer by rothauserc(4718)

Note that we do not need the 6mm measurement for surface area calculations!
Surface Area (SA) of cylinder (two ends are circles) formula is
SA = 2*pi*radius squared + 2*pi*radius*height
SA = 2*22/7*4 squared + 2*22/7*4*10, therefore SA = 100.6 + 251.4 = 352 square mm's
Surface area of cone (one end is a circle) formula is
SA = pi*radius squared + pi * radius * hypotenuse (s)
first calculate s squared = radius squared + height squared,
s squared = 10 squared + 25 squared = 100+625=725
s = square root of 725 = 26.9 mm
SA = 22/7*10 squared + 22/7*10*26.9 = 2200/7 + 22/7*10*26.9=314.3+825.4=1159.7 square mm's
Surface area of rectangular solid is
the square base is 40mm by 40mm = 1600 square mm and there are two ends so total for 2 square ends is 3200 square mm
now there are 4 sides that are 50 by 40 or 2000 square mm * 4 = 8000 square mm
adding the two calculations
SA = 3200 + 8000 = 11200 square mm's